Aureo:

Convert to vertex form. f(x)=x^2+8x+13

KylieAurealis:

y=2(x-2)2+21

KylieAurealis:

The 3rd 2 is supposed to be to the second

dude:

Vertex form is written in \(y=a(x-h)+k\) In which the vertex is (h,k)

KylieAurealis:

Nvm XD i suck at this

dude:

To find the vertex use the equation \(\large x=\frac{-b}{2a}\)

KylieAurealis:

I wouldnt know. Im in geometry lol

Aureo:

So it'd be \[\frac{ -8 }{ 2 }\] Which is -4

dude:

Yes so that is the x value, now substitute that into the original equation to find y \(y=(-4)^2+8(-4)+13\)

Aureo:

Okay so that would be y = -3?

dude:

Right that is the vertex (-4,-3)=> (h,k) So now we just substitute that into \(y=a(x-h)+k\) =>\(y=(x-(-4))-3\) or \(\boxed{y=(x+4)^2-3}\)

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude Vertex form is written in \(y=a(x-h)+k\) In which the vertex is (h,k) \(\color{#0cbb34}{\text{End of Quote}}\) Correction : \(y=a(x-h)\color{red}{^2}+k\)

Aureo:

Thank ye kind sir.